OpenStudy (anonymous):

what is (6.5*10^-7)(4*10^-10)

OpenStudy (mr.coolio):

2.6 x 10^-16

OpenStudy (anonymous):

10^-16 is not one of the awnsers

OpenStudy (mr.coolio):

10 to the power of -16 is not among your answers

OpenStudy (anonymous):

how r u changeing your picture

OpenStudy (anonymous):

and no

OpenStudy (mr.coolio):

oh i went into settings and changed it just now

OpenStudy (mr.coolio):

oh sry i do not know the answer then

OpenStudy (mr.coolio):

I used this website to figure out the answer and it did not change i do not know sry http://www.wolframalpha.com/input/?i=%286.5*10%5E-7%29%284*10%5E-10%29

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

HELP

OpenStudy (tkhunny):

Who cares if 10^(-16) is not one of the answers? Find the one that is! (6.5*10^-7)(4*10^-10) = 6.5 * 4 * 10^(-17) = 26 * E(-17) = 2.6 E(-16) = 260 E(-18) Don't just punt. Go hunting!